OpenStudy (anonymous):

A bowl contains candies of the same size in three flavors: orange, strawberry, and pineapple. If the probability of randomly pulling out an orange candy is 1 over 9, and the probability of randomly pulling out a strawberry candy is 3 over 9, what is the probability of randomly pulling out a pineapple candy?

OpenStudy (anonymous):

so would it be 4/9 ? cause i did 1 + 3 and kept the denominator the same.

OpenStudy (anonymous):

The sum of all the probabilities must be 1. What do think the answer is?

OpenStudy (anonymous):

why would it be 5/9 ? @pinkbubbles

OpenStudy (anonymous):

P(orange) + P(strawberry) + P(pineapple) = 1 1/9 + 3/9 + P(pineapple) = 1 Understand?

OpenStudy (anonymous):

i think so. so would my answer then be 2/9?

OpenStudy (anonymous):

Not quite.\[\frac{ 1 }{ 9 } + \frac{ 3 }{ 9 } + P(pineapple) = \frac{ 9 }{ 9 }\]What is P(pineapple)?

OpenStudy (anonymous):

oh I think I got it now, so pineapple is 5/9 right?

OpenStudy (anonymous):

Bingo!! Well done.

OpenStudy (anonymous):

thank you, I really understand it now! :)

OpenStudy (anonymous):

That's the most important thing. Keep up the good work.